Gene Frequencies We will let p designate the frequency (=proportion) of gene B in a population, and we will let q designate tPlease show how to works with type. Thank you.
Answer #1

Answer: Hardy-Weinberg principle is explained in the give example.

Before applying the principle to a population, following conditions need to be satisfied.

· Population should be large

· There should be random mating between the individuals of the species.

· No genetic drift taking place because of migration, mutation etc.

· No selection of a particular allele

If all these conditions are satisfied then the population is said to be in H-W equilibrium, according to this principle the allele frequencies in the population will remain the same over the generations

If we consider any character and if there are two alleles for the character

Let’s consider the alleles as B and b

If B allele represents the dominant character then it will expressed in homozygous BB as well as heterozygous condition Bb

And the recessive allele will express itself only in homozygous condition i.e. bb

The allele frequencies are given by the equation

p2+2pq+q2 =1   i.e. p+q =1

p2 = BB= frequency of the homozygous dominant genotype.

2pq = Bb= frequency of heterozygotes in the population.

q2 = bb= frequency of the homozygous recessive genotype

p = frequency of the dominant allele, note that dominant allele B will appear in homozygous as well as in heterozygous condition.

q = frequency of the recessive allele, note that recessive allele b will appear in homozygous as well as in heterozygous condition.

Now to start with the calculation, recessive character is considered first as it will appear only in homozygous condition.

If the question is asked, make sure what is given in the question, is it the frequency of the individuals with the particular character or the frequency of the allele is mentioned. Accordingly fill the values of p and q

In the given example they have given the frequency of the recessive allele that means q is given

q= 0.3 therefore the frequency of the homozygous recessive individuals will be

q2 that will be equal to 0.3×0.3 =0.09, if you multiply this with 100, you will get the percentage of individuals in the population who are showing the recessive character (recessive phenotype)

Now same way value of p can be calculated by using the formula p+q=1

Therefore p=1-0.3 that is p=0.7 this is the frequency of B allele in the population.

To calculate the homozygous dominant individuals BB, you need to calculate p2 that will be equal to 0.7×0.7 that is 0.49. Hence the percentage of homozygous individuals in the population is 49%

To calculate the heterozygous population, you need to consider 2pq that will be equal to 2×0.7×0.3=0.42, hence the percentage of heterozygous individuals in the population is 42%

If you add the both the above percentages (as dominant character is expressed in homozygous as well as heterozygous condition)

it will be equal to 91%, the same value which we get if consider the percentage of recessive individuals as 9% which we got from the first calculation.

To summarise

In the given population, if the phenotypes are considered then the number of individuals expressing the dominant phenotype = 91%

And the number of individuals expressing the recessive phenotype =9%

If the genotypes are considered then the frequency of B allele in the population is 0.7

And frequency of b allele in the population is 0.3.

You can cross check by putting the values in the equation p2+2pq+q2 =1  

Note: suppose in the problem, the number of individuals are given then you have to consider p2 or q2 depending on whether dominant or recessive trait is considered as in the following example, only the first line of the problem I have mentioned

The number of individuals in the population expressing the recessive character is 36%

By reading this it can be concluded that q2 is given which is 0.36 therefore p2 in this case will be 64% that is 0.64

Now the further calculations can be done with this information.